OpenStudy (anonymous):

Please help! Find Tan x/2, given that tan x=-3 and x terminates in 90 degrees < x <180 degrees.

OpenStudy (anonymous):

@calmat01 - can you help with this one?

OpenStudy (anonymous):

This is a half angle problem. We need to use a half angle formula for it.

OpenStudy (anonymous):

ok. the square root one?

OpenStudy (anonymous):

That would be one we could use. But perhaps we could use eithr of the ones without the square root.

OpenStudy (anonymous):

the one with the sin x in the numerator or denominator?

OpenStudy (anonymous):

Let's use the one with the sine in the denominator.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok, first let us draw a triangle that fits the description of your problem.

OpenStudy (anonymous):



OpenStudy (anonymous):

Awesome! I wish I had 25 of you in my classroom!

OpenStudy (anonymous):

you just made my daughter feel great! she took the reigns and drew it. She is feeling much better about this trig now than a couple of hours ago. You are fabulous!!!!

OpenStudy (anonymous):

I have been doing this for 20 years. I know a good student when I see one. All it takes is a little guidance, and practice. She's the type of student I wish I had in all of my classes.

OpenStudy (anonymous):

Ok, so, now we have our triangle, just fill in your values for cosine and sine of x.

OpenStudy (anonymous):

\[(1-(- \sqrt{10}) /10)/((3\sqrt{10})/10)\]

OpenStudy (anonymous):

Good job. Now simplify it. BTW, you wouldn't have had to rationalize the denominator in this case because they will just cancel out in the end.

OpenStudy (anonymous):

i dont know how to simplify this

OpenStudy (anonymous):

Think of it as a fraction divided by another fraction. Our rule is to multiply by the reciprocal. So, it becomes\[\frac{ 1+\sqrt{10} }{ 10 } \times \frac{ 10 }{ 3\sqrt{10} }\]